OpenStudy (ujjwal):

If |3+4i-Z|=5 then the locus of complex number Z is

OpenStudy (anonymous):

Put z = a + ib Then find the magnitude on left hand side..

OpenStudy (anonymous):

Firstly you will have to combine like terms and then take the magnitude..

OpenStudy (apoorvk):

Hmm, the locus of such a format, i.e. \(|z - z_1| = a\) represents a circle.

OpenStudy (anonymous):

Yes but I will show you how..

OpenStudy (ujjwal):

yeah, i did that! And i get a circle. But then i even think it's a hyperbola.. So, which one is correct?

OpenStudy (anonymous):

Let us solve it and then we can say what it is..

OpenStudy (ujjwal):

Ok!

OpenStudy (anonymous):

u kno it's a circle :|

OpenStudy (anonymous):

On combining we get the magnitude as: \[\sqrt{(3-a)^2 + (4-b)^2} = 5\] Yeah it represents the Circle..

OpenStudy (apoorvk):

it IS a circle. let me show you how. If the real-axis is the X-axis, and the imaginary axis is the Y-axis, and \(z = x+iy\), then: \[| 3 + 4 \iota - x + \iota y\ | = 5\] \[or, ~ \sqrt{(3-x)^2 + (4+y)^2} = 5\] Now when you square both sides, \[or, ~ (3-x)^2 + (4+y)^2 = 5^2\] And this, represents a circle!

OpenStudy (anonymous):

Where (3,4) is the centre and radius is 5..

OpenStudy (apoorvk):

Aye^

OpenStudy (ujjwal):

@apoorvk , Sth is wrong in your reply! that should be: |3+4i-x-iy|=5

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

@apoorvk centre will be(3,4) and not (3, -4)..

OpenStudy (anonymous):

@ujjwal you are saying right..

OpenStudy (apoorvk):

Oh yeah am sorry! Please accommodate for that! thanks @waterineyes Should be: \[|3+4ι−x-ιy |=5\] \[or, \sqrt{(3−x)^2+(4-y)^2}=5\] Now when you square both sides, \[or,~ (3−x)^2+(4-y)^2=5^2\]

OpenStudy (apoorvk):

Thanks @ujjwal as well

OpenStudy (anonymous):

Or you can write it as: \[(a-3)^2 + (b - 4)^2 = 5^2\] Compare with the Standard Equation: \[(x-h)^2 + (y-k)^2 = r^2\] So you will get (h,k) = (3,4) and Radius = 5..

OpenStudy (ujjwal):

Shouldn't the magnitude be: \[\sqrt{(3-x)^2+i^2(4-y)}=\sqrt{(3-x)^2-(4-y)}\]

OpenStudy (ujjwal):

*(4-y)^2

OpenStudy (ujjwal):

I am concerned about - sign

OpenStudy (anonymous):

See, if you have complex number as : a + ib Then magnitude is given by: \[r = \sqrt{a ^2 + b^2}\]

OpenStudy (apoorvk):

Nope - when I take the magnitude, I do not need the 'iota' thingy - because magnitude physically represents the distance from the origin in the Argand's plane, and iota is like a 'direction-provider' here.

OpenStudy (anonymous):

i is not considered in magnitude...

OpenStudy (ujjwal):

Thanks @apoorvk and @waterineyes !! Thanks a lot!

OpenStudy (anonymous):

Welcome dear..

OpenStudy (apoorvk):

Anytime! ;]